OpenStudy (anonymous):

Please help will medal! e ^a−9 +1=34.2

OpenStudy (anonymous):

does the -9 belong with the 'a' or is it e^(a) - 9 + 1 = 34.2

OpenStudy (anonymous):

that's the equation

OpenStudy (anonymous):

ok so you can turn it into e^(a-9)=33.2 Remember what I said about the natural log ln? e^q = x <=> q= ln(x) That would make a-9=ln(33.2) Calculate ln(33.2) in your calculator and then a-9= whatever number you got from your calculator. a= whatever number you got + 9

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No prob :-)